# ConTextTab: A Semantics-Aware Tabular In-Context Learner

### v2

@@ 4.2 Evaluation
Datasets:  We use a variety of tabular prediction datasets to evaluate and compare our approach to established baselines and other SOTA methods. Namely, we evaluate all models on the following benchmarks: OpenML-CC18 [1], a pure classification benchmark; OpenML-CTR23 [10], a pure regression benchmark; TALENT [37], a recently introduced diverse benchmark containing over 300 classification and regression benchmarks. Here, we focus on a subset containing 45 datasets that are representative of the overall performance of the baselines investigated in the original works, which we refer to as the TALENT-Tiny benchmark; TabReD [30], a small but challenging benchmark of large datasets representative of practical prediction tasks; and finally CARTE [22], a mixed classification and regression benchmark containing highly semantic features and few numerical ones.
@@
Across all benchmarks, we evaluate 91 regression and 112 classification tasks, ranging from 5 to 3 k columns and 400 to roughly 400 k training examples. Due to the large number of evaluated datasets, we do not perform cross-validation but evaluate a fixed test split for each task. For models that do not explicitly use a validation split, we concatenate the train and validation splits. We refer to Appendix B.2 for a visualization of the dataset statistics and further details.
@@
Baselines:  We compare our approach to several established classical methods as well as recent ICL and other deep learning models. In particular, we compare with TabPFN [17], TabICL [28], and TabDPT [24] as the most recent table-native ICL methods. For all pretrained models, we use the latest available release and checkpoints as of May 2025. As the SOTA on the CARTE benchmark, we also evaluate the CARTE model proposed by [22]. As another well-performing recent deep learning approach, we evaluate RealMLP [18]. Note that both CARTE and RealMLP need to be finetuned or fitted for each dataset separately. We perform per-dataset hyperparameter tuning via random search for RealMLP as provided in the author’s implementation via the pytabkit package. Also using the implementation wrapper from pytabkit, we evaluate SOTA boosted tree baselines, namely XGBoost [2], LightBGM [21], and CatBoost [27] – both in the meta-tuned default variants (TD) [18] as well as via a Parzen-tree estimator-based hyperparameter optimization (HPO). Furthermore, we evaluate several common, non-tuned baselines from scikit-learn [25], namely naive, linear, and KNN estimators, as well as a random forest and histogram-based gradient boosting tree estimators. Finally, as the gold standard in tabular prediction, we evaluate AutoGluon [9], an AutoML solution stacking and ensembling many of the baselines outlined above. Please refer to Appendix B.1 for full details on the used baseline versions and parameters as well as details about the dataset preprocessing and encoding.
@@
Metrics:  Throughout, we show (mean) accuracy for classification tasks as well as (mean) $R^{2}$ score for regression tasks. As averaging across a large number of datasets with varying performance can blur relative performance across models, we also evaluate (mean) rank. Within each benchmark, we calculate the mean rank across all constituent datasets and models. Full rank performance is averaged across all datasets across all benchmarks, not as per-benchmark averages as some benchmarks contain only few tables. To reduce the impact of small noise, we count model performances as ties when their evaluation scores lie within 0.005 of each other.
@@ 5 Results
Table 1: Performance comparison across all evaluated benchmarks, depicting mean accuracy (Acc) for classification and $R^{2}$ score for regression tasks, in percent. Missing values, due to architectural limitations or failed evaluations, are denoted as N/A and excluded from the mean rank calculation.
Model Name All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 AutoGluon 78.7 73.8 88.5 67.0 86.0 64.6 87.9 73.7 ConTextTab 2.20 1.49 76.9 72.2 2.79 86.8 3.29 72.9 3.88 85.4 63.4 2.30 87.7 76.1 TabPFN 2.66 5.08 72.4 65.0 2.67 87.0 2.63 74.9 3.00 85.6 63.8 1.73 87.3 75.1 LightGBM [HPO] 2.85 4.06 72.8 66.1 3.26 86.9 3.31 61.9 1.62 85.9 64.4 3.19 86.4 72.4 XGBoost [HPO] 3.16 4.47 72.6 65.7 4.00 86.3 3.17 71.4 1.75 86.0 64.2 3.24 86.2 72.6 CatBoost [HPO] 3.24 4.02 75.6 66.1 3.57 86.7 4.40 -44.4 2.00 85.8 63.7 3.76 86.0 72.1 HistGradBoost 3.46 5.08 72.5 64.8 3.62 86.1 4.80 51.2 1.75 85.9 63.9 3.62 86.3 67.6 RealMLP [TD] 3.91 8.39 70.2 59.6 3.18 87.2 3.80 45.7 1.25 86.0 64.6 4.08 86.2 71.5 Random forest 4.56 7.08 71.5 63.3 4.69 85.7 5.17 53.6 5.38 85.4 60.7 4.54 85.8 70.6 Naive 8.61 11.22 53.0 -1.8 9.39 47.0 9.69 -8.5 8.62 80.8 -0.6 9.97 53.4 -22.3 CARTE N/A 3.16 76.1 68.5 N/A N/A N/A N/A N/A N/A N/A N/A N/A N/A TabDPT N/A 5.02 72.7 65.1 2.71 87.8 2.46 72.4 N/A N/A 60.9 2.65 86.7 74.8 TabICL N/A N/A 72.5 N/A N/A 88.0 N/A N/A N/A 85.1 N/A N/A 87.4 N/A
@@
The main results, evaluating ConTextTab and the baselines as outlined in the previous section, are shown in Table 1. Here, we exclude AutoGluon from the overall rank and best-model comparison as it is ensembling and stacking a multitude of model architectures, making it difficult to highlight architectural strengths and weaknesses. Overall, our model achieves the highest rank across all evaluated datasets. This performance gain is statistically significant, as depicted in the critical difference diagram in Figure 2 (left). Our model performs particularly well for the semantically rich CARTE benchmark, highlighting the importance of incorporating semantic understanding into tabular ICL. Note that TabPFN, not incorporating semantics, performs worse than both tuned and non-tuned boosted trees in this particular instance. On the other end, while the absolute performance of ConTextTab for large datasets, e.g. within the TabReD benchmark, is not bad, it falls behind tuned boosted trees here, which still achieve SOTA, on par with AutoGluon.
@@
To investigate the dependence on dataset size more closely, we show the performance on the CARTE benchmark across varying subsampled sizes, ranging from 128 rows to the full dataset, in Figure 3. ConTextTab consistently outperforms other models across all sample sizes and even surpasses AutoGluon for up to 2048 training samples. This highlights the strong capabilities of tabular ICL but also the need for further research into scaling these architectures to effectively deal with much larger context sizes as well as training datasets. We investigate the dataset-size dependent rank performance across all 203 datasets in Appendix A.1, further highlighting these challenges.
Additional results, including additional baselines and win ratios, are given in Appendix A.2.
Figure 2: Left: critical difference diagram between ConTextTab and several baselines, across all 203 evaluated datasets. Right: Impact of pretraining dataset size on validation accuracy and $R^{2}$ scores.
Figure 3: Average rank, accuracy, and regression results on the CARTE benchmark across various data subsets, ranging from 128 rows to the full size.
5.1 Ablation studies
In the following, we discuss several experiments ablating aspects of our model reported in Table 2.
Table 2: Multiple ablations as relative differences with respect to one of two base configurations, using binning (top) or one-dimensional embedding (bottom) for regression. Improvements over the base model are highlighted in bold.
Experiment All CARTE OML-CC18 OML-CTR23 TabReD TALENT-Tiny Rank Rank Acc R2 Rank Acc Rank R2 Rank Acc R2 Rank Acc R2 base model (binning) 2.28 1.76 76.0 71.4 2.57 86.8 2.66 70.9 1.00 85.4 62.9 2.32 87.2 73.5 large size 2.76 2.18 +0.2 -0.2 4.01 -5.3 1.77 -0.7 3.75 -25.4 -13.5 1.84 -0.1 -0.0 medium size 5.44 8.45 -5.6 -3.7 4.76 -1.2 3.49 -1.0 5.75 -9.3 -3.5 4.41 -2.6 -2.4 small size 7.68 9.98 -2.9 -6.8 6.57 -3.0 8.17 -4.5 8.12 -1.7 -3.8 6.14 -2.7 -6.7 mini size 8.10 10.69 -4.2 -8.4 7.15 -4.0 8.06 -4.2 5.50 -0.3 -3.5 7.00 -3.7 -6.9 multilingual-e5-small 3.22 4.16 -0.3 -1.5 3.35 -0.4 2.77 -0.1 2.38 -0.2 -1.3 2.27 +0.7 -0.7 gte-multilingual-base 3.22 3.59 +0.0 -1.2 3.18 -0.3 3.60 -1.4 1.62 -0.2 -0.9 2.76 +0.4 -1.3 curriculum learning 2.28 1.51 +1.1 +0.1 3.10 -0.2 1.80 -0.1 1.38 +0.4 -0.2 2.38 -1.0 +0.4 clustering N/A N/A -0.1 N/A 3.82 -1.5 N/A N/A N/A -0.6 N/A N/A -1.3 N/A w/o bagging 2.82 2.76 -0.4 -0.4 3.26 -0.7 3.03 -0.4 1.00 +0.1 -0.5 2.24 -0.6 -0.0 ISAB 5.48 6.16 +0.3 -3.1 5.57 -2.1 5.51 -0.9 3.88 -0.2 -2.7 4.70 -0.5 -6.1 non-shared weights 4.36 5.61 -0.1 -2.6 4.29 -0.9 3.09 +0.8 5.88 -1.0 -2.6 3.65 -0.7 -1.2 base model (1-dim) 2.00 1.80 76.7 72.0 2.36 86.9 2.06 65.0 1.75 85.3 63.3 1.54 87.9 75.8 0.1% clipping 2.03 1.80 0.0 0.0 2.36 0.0 2.37 -0.3 1.38 0.0 +0.2 1.54 0.0 0.0 0.5% clipping 2.08 1.76 0.0 0.0 2.32 -0.1 2.60 +4.0 1.38 +0.1 +0.1 1.70 +0.1 -0.2 2% clipping 2.97 2.88 0.0 -0.5 2.43 -0.1 5.03 +5.8 1.75 +0.1 -0.5 2.43 0.0 -2.0 binning 4.04 4.69 -0.7 -0.6 3.21 -0.1 6.34 +5.9 1.12 +0.1 -0.3 3.24 -0.7 -2.3 context=20000 2.06 1.39 +0.3 +0.6 2.62 -0.2 1.66 +4.2 3.25 -0.4 -0.5 2.00 -0.4 +0.5 context=10000 2.00 1.61 +0.3 +0.3 2.64 -0.2 1.57 +4.2 1.75 0.0 -0.1 1.76 -0.2 +0.5 context=4096 2.81 3.55 -0.1 -0.6 3.03 -0.3 2.11 +4.0 2.25 -0.3 -0.1 2.16 -0.4 +0.4 context=2048 4.11 6.41 -1.0 -1.9 3.61 -0.4 3.06 +3.6 2.50 -0.1 -0.7 3.24 -0.9 +0.1 context=1024 5.49 9.10 -1.3 -3.8 4.33 -0.5 4.60 +2.7 2.62 -0.4 -1.6 4.24 -1.3 -0.6 fine-tuning 2.60 2.51 +0.3 +0.1 3.1 +0.2 1.94 +7.2 2.62 +0.2 -0.3 2.38 -0.1 +0.8
Model size:  As expected, decreasing from “base” to smaller models impacts performance. On the other hand, increasing to “large”, thus doubling the number of layers and increasing the hidden dimension, has no statistically significant impact on performance. As observed in Figure 2 (right), increasing the amount of training data leads to stagnating validation performance gains. We hypothesize that, in the current setup, our model is likely limited by the amount or diversity of data available with T4.
Context size:  The trained models can scale the input context size at inference. As expected, longer context improves results in a monotonic way, in exchange for longer computation time. In practice, the effects are stronger on larger datasets, simply because shorter context models (even with bagging) only get to see a subset of the training data. We stress again that this is entirely extrapolating outside of the pretraining setup, where the largest observed context has size 950. Nevertheless, with larger datasets and context, the model performance saturates, requiring further research in how to increase the capacity of tabular ICL.
Training data:  We ran training on randomly chosen subsets of the T4 dataset of $10^{n}$ tables, $n\in\{1,\dots,6\}$. For this experiment we report the validation scores (on the validation held-out subset of the CARTE dataset) in Figure 2 (right). The results suggest that at least 100 k of the tables are needed in order to train a model of state-of-the-art quality. While exact numbers may vary when using other datasets or model sizes, this gives a high-level estimate of the amount of required data, giving indication of why our model could not profit from a second step of curriculum learning on the much smaller dataset used by Ma et al. [24]. Potentially, training is still limited by data amount or diversity, e.g. missing longer and wider tables, likely required to further scale tabular ICL.
Classification target:  We compare the default cross-entropy loss for classification with the clustering approach described in Section 3.4, allowing to retain the semantics of the target labels, lifting the limitation on the number of target classes as present in TabPFN or TabICL. However, most evaluation datasets only contain non-descriptive labels, and the number of classes is limited. Hence, the potential of clustering cannot emerge, and the default maintains a modest edge. We believe that semantically richer high-cardinality benchmarks are required to spur research in this direction.
ISAB block:  We observed comparable results on classification tasks and a modest drop in performance on regression tasks, while substantially enhancing efficiency by eliminating quadratic time complexity and reducing runtime by a factor of ten for large contexts.
Bagging:  As expected, using bagging at inference time leads to a consistent improvement of results. At the same time, in absolute terms, the gap is rather marginal, probably thanks to the fact that ConTextTab gets rid of most sources of non-equivariance in the architecture (e.g. category indexing and positional encoding). As a result, depending on the use case, this allows to drop bagging, or reduce the number of bags, without compromising performance significantly.
(No) weight sharing:  We observed a slight performance degradation, which may be due to either an insufficient training duration or suboptimal selection of training hyperparameters.
Further ablations:  Other changes (the used sentence embedding model, longer context in training, target clipping for regression, as well as benchmark-dependent fine-tuning) did not result in significant performance gains. Therefore, we refer to the Appendix A.3 for further discussion of these ablations.
6 Conclusions
Limitations and future work:  While achieving SOTA results across the investigated datasets, we observe several limitations of our proposed approach. One drawback of using real-world data for training is the possibility of contamination, e.g. the presence of evaluation tasks in the training corpus. Since the CARTE benchmark is our focus, we conducted a contamination study, using the column name and cell value embeddings created by a text embedder and matching similarities. We did not find any contamination of CARTE in T4. For contamination of OpenML datasets, we refer to the original study by [11]. Either way, as a single table is only seen a few times during training, we believe that memorization is likely not a practical problem when training on real-world data.
Generally, all investigated table-native ICL models, e.g. ours, TabPFN, TabDPT, and TabICL, fail to scale their performance to very large datasets. Increasing the context length did not resolve these issues. Using local context, as done in TabDPT, might overcome such limitations but we were unable to evaluate TabDPT on TabReD as it could not handle large tables. In the large-data case, conventional methods, in particular when stacked and ensembled via AutoGluon, still perform best. Overcoming this remains one of the major research challenges for tabular foundation models.
We observe little gain when training on larger data corpora, e.g. increasing the pretraining subset of T4 or adding tables from the AutoML benchmark. Together with the investigated scaling of our model, we believe that more diverse real-world data is needed to fully unlock semantic understanding at scale. While models trained on synthetic data do not face this issue in principle, how to incorporate semantic alignment in these models remains an open question. Similarly for evaluation, semantically rich benchmarks are scarce. While CARTE includes rich semantic features, all classification targets are binary, containing no semantic information. Furthermore, it contains mostly small to medium sized datasets, in particular in terms of the number of features. Hence, we argue that more diverse, semantically rich data is required – both at scale for pretraining larger models with longer context, as well as curated for evaluation.
Finally, our investigation to utilize target semantics and lift the strict class limit at inference prevalent in current tabular ICL, by using a dynamic supervised clustering approach, did not result in performance gains. This might relate to the lack of semantically meaningful targets in current tabular benchmarks as well as the limited target cardinality across them.
Summary:  We have introduced ConTextTab – a context aware table-native in-context learner trained on real-world data for both classification and regression tasks. Evaluated across a wide range of tabular benchmarks – including both classification and regression tasks – our model achieves state-of-the-art results significantly outperforming existing table-native ICL approaches such as TabPFN, in particular on the semantically rich CARTE benchmark and low-data regime. We outline future research efforts required to further improve semantic understanding in tabular foundation models, in particular the need for more diverse semantically rich pretraining as well as evaluation data.
Acknowledgements
We would like to thank Johannes Hoehne, Johannes Hoffart, and Markus Kohler for their insightful comments and suggestions throughout the development of this work, which have greatly contributed to shaping its direction and quality.
Appendix A Further Results
A.1 Relation between dataset size and model performance
We plot the average rank of each model as a function of the dataset size (expressed in number of rows) across all evaluated datasets in Figure 4. Note that TabICL is missing from this comparison as it only handles classification tasks. We can observe that, as expected, ConTextTab and TabPFN excel in the low data regime, while AutoGluon takes the lead when enough data is available. In the very low data regime below 1000 training rows, TabPFN performs best, however our model surpasses its performance for larger datasets with more than 1000 rows. Overall, ConTextTab remains competitive with AutoGluon until roughly 10 000 rows, possibly as a result of the limitation in both the inference context size and the context size seen during training. After 10 000 rows, gradient boosting methods, as well as RealMLP, start surpassing the performance of ConTextTab as well as TabPFN. Overall, this indicates the need of further research for tabular ICL to handle larger amounts of training data, but also the availability of more diverse benchmarks, covering larger datasets, as the current evaluation is dominated by datasets with less than 10 k rows and less than 100 columns, see Figure 6.
Figure 4: Relation between number of training dataset rows and performance, obtained as a LOWESS regression in the plane $\log(n_{\mathrm{rows}},\textnormal{rank})$. The confidence bands are the 80% confidence intervals obtained via bootstrapping.
A.2 Extended results
Including further baselines not depicted in the main paper, Table 3 shows all collected evaluations across the investigated benchmarks. Furthermore, we show the win ratio confusion matrix and averages of the investigated models in Figure 5.
(a) Win ratio matrix with Model A wins over Model B.
(b) Average Win ratio of Model A against all others.
Figure 5: Win ratio confusion matrix and average of the investigated models across all 203 datasets. Wins are calculated based on accuracy on classification and $R^{2}$ on regression datasets. Ties are not counted as wins. Models are sorted by descending overall rank. Note that CARTE, TabDPT, RealMLP [HPO], and TabICL were not successfully evaluated on all datasets and only compared on evaluated ones, resulting in a smaller support and skewing their win rates. Therefore, they are separately grouped last, not used in averaging, and only shown for completeness.

### v4

@@ 4.2 Evaluation
Datasets:  We use a variety of tabular prediction datasets to evaluate and compare our approach to established baselines and other SOTA methods. Namely, we evaluate all models on the following benchmarks: OpenML-CC18 [openmlcc18], a pure classification benchmark; OpenML-CTR23 [openmlctr23], a pure regression benchmark; TALENT [talent], a recently introduced diverse benchmark containing over 300 classification and regression benchmarks. Here, we focus on a subset containing 45 datasets that are representative of the overall performance of the baselines investigated in the original works, which we refer to as the TALENT-Tiny benchmark; TabReD [tabred], a small but challenging benchmark of large datasets representative of practical prediction tasks; and finally CARTE [CARTE], a mixed classification and regression benchmark containing highly semantic features and few numerical ones.
@@
Across all benchmarks, we evaluate 91 regression and 112 classification tasks, ranging from 5 to 3 k columns and 400 to roughly 400 k training examples. Due to the large number of evaluated datasets, we do not perform cross-validation but evaluate a fixed test split for each task. For models that do not explicitly use a validation split, we concatenate the train and validation splits. We refer to Appendix LABEL:app:dataset for a visualization of the dataset statistics and further details.
@@
Baselines:  We compare our approach to several established classical methods as well as recent ICL and other deep learning models. In particular, we compare with TabPFN [tabpfnv2], TabICL [tabicl], and TabDPT [tabdpt] as the most recent table-native ICL methods. For all pretrained models, we use the latest available release and checkpoints as of July 2025. As the SOTA on the CARTE benchmark, we also evaluate the CARTE [CARTE]. As additional well-performing recent deep learning approaches, we evaluate RealMLP [realmlp] and TabM [tabm]. Furthermore, we evaluate SOTA boosted tree baselines, namely XGBoost [xgboost], LightBGM [lightgbm], and CatBoost [catboost]. The tree baselines as well as RealMLP and TabM are evaluated in the default (D) or meta-tuned default variants (TD) [realmlp], as well as via Parzen-tree estimator-based hyperparameter optimization, either with a fixed holdout set (HPO) or ensembled via 5-fold inner cross-validation with per-split HPO (HPO, CV). For this, we use the implementation provided by pytabkit [realmlp]. Additionally, we evaluate several common, non-tuned baselines from scikit-learn [sklearn], namely naive, linear, and KNN estimators, as well as a random forest and histogram-based gradient boosting tree estimators. Finally, as the gold standard in tabular prediction, we evaluate AutoGluon [autogluon], an AutoML solution stacking and ensembling many of the baselines outlined above. We refer to Appendix LABEL:app:baselines for full details on the used baseline versions and parameters as well as details about data preprocessing and encoding.
Metrics:  Throughout, we show (mean) accuracy for classification tasks as well as (mean) $R^{2}$ score for regression tasks. Since we have observed some severe outliers in some regression tasks that greatly affected average scores, we perform a soft-clipping. That is, all negative $R^{2}$ scores are mapped to [1,0)[-1,0) via $\tanh$ while keeping positive scores untouched. This retains the relative ordering of scores while being smoothly differentiable and less distorting to the mean.
@@
As averaging across a large number of datasets with varying performance can blur relative performance across models, we also evaluate (mean) rank. Within each benchmark, we calculate the mean rank across all constituent datasets and models. Full rank performance is averaged across all datasets across all benchmarks, not as per-benchmark averages as some benchmarks contain only few tables. To reduce the impact of small noise, we count model performances as ties when their evaluation scores lie within 0.005 of each other. Furthermore, we report critical difference (CD) diagrams, win ratio matrices, and p-values enabling 1 vs 1 model comparisons in the Appendix.
@@ 5 Results
The main results, evaluating ConTextTab and baselines, are shown in Table 5. Here, we exclude AutoGluon from the overall rank and best-model comparison as it is ensembling and stacking a multitude of model architectures, making it difficult to highlight architectural strengths and weaknesses. Overall, our model performs competitive across all non-semantic benchmarks while setting a new standard on the semantically rich CARTE benchmark, where it ranks best. In fact, this is statistically significant over all models but CatBoost and AutoGluon, both of which are extensively tuned and ensembled, as shown in the critical difference diagram for CARTE in Figure LABEL:fig:critical_difference_and_pretraining_size (left). Notably, ConTextTab is significantly better than other tabular ICL approaches, which highlights the importance of incorporating semantic understanding. On the other hand, TabPFN, not incorporating semantics, performs worse than tuned as well as untuned trees on CARTE. For details, we refer to Appendix LABEL:app:additional-results.
On the other end, the absolute performance of ConTextTab for non-semantic benchmarks, while good, falls behind tuned and ensembled boosted trees as well as RealMLP, which achieves SOTA in terms of ranking across all evaluated datasets (albeit at a much higher training time and cost). Nevertheless, these differences in ranking are mostly not statistically significant. To this end we report per-benchmark CD diagrams as well as win-ratio and p-values for detailed model comparison in Appendix LABEL:app:additional-results-extended. Namely, ConTextTab does not perform significantly worse than the best-ranking model on all non-semantic benchmarks excluding OpenML-CTR23 (where it is however not significantly different to tuned ensembled trees).
Hence, ConTextTab sets a new standard on semantically rich datasets while performing on-par to existing approaches on non-semantic ones. Nevertheless, the performance of AutoGluon generally indicates further headroom to improve single-model performance and tabular ICL in particular.
@@
Finally, tuned trees perform particularly well for large datasets as represented by the TabReD benchmark, even outperforming AutoGluon in some instances. Motivated by this, we investigate the dependence on dataset size more closely and the performance on the CARTE benchmark across varying subsampled sizes, ranging from 128 rows to the full dataset, in Figure LABEL:fig:carte-sub. ConTextTab consistently outperforms other models across all sample sizes and even surpasses AutoGluon for up to 2048 training samples. This highlights the strong capabilities of tabular ICL but also the need for further research into scaling these architectures to effectively deal with much larger context sizes as well as training datasets. We investigate the dataset-size dependent rank performance across all 203 datasets in Appendix LABEL:app:additional-results-data-size, further highlighting these challenges.
Additional results, including additional baselines, CD diagrams, as well as 1 vs 1 win ratios and p-values, are provided in Appendix LABEL:app:additional-results-extended. Furthermore, we refer to Appendix LABEL:app:runtime-analysis for an analysis of the model runtime in comparison to other baselines.